Use linear combinations to slove the linear system. Then check your solution.
3g-24=-4h
-2+2h=g
Below is what I have tried, however it was unsuccsessful:
3g+4h=-24
-g-2h=2
_________
-3/1g (fraction)=3 =-21
_________________________
3g+4h=-24-------3g+4h=-24
-3(-g-2h=2)-----3g-2h=2
_________________________
6h=-21/6=-3.5



3(3.5)-4h=-24
9.5-4h=-24
= 9.5
_____________
-33.5/4 (divide)= -8.25

Answer by stanbon(75887)

Use linear combinations to slove the linear system. Then check your solution.
3g-24=-4h
-2+2h=g
--------------
Rearrange the equations:
3g + 4h = 24
g - 2h = -2
---------------
Multiply the 2nd equation by 2 to get:
2g -4h = -4
--------------
Add that equation to the 1st to get:
5g = 20
g = 4
--------------
Substitute that into 3g+4h = 24 to solve for "h":
3*4 + 4h = 24
4h = 12
h = 3
----------------
Solution: h=3 and g = 4
============================
Cheers,
Stan H.